Construct a quadratic equation in x with solutions 4 and -2 (steps would be appreciated if possible)
Answer by london maths tutor(243)

x = 4
x-4 = 0
x= -2
x + 2 = 0
Therefore:
(x-4)(x+2) = 0
x^2 + 2x - 4x - 8 = 0
x^2 - 2x - 8 = 0